greatest common factor of 63, 147, and 273
Answer by Earlsdon(6294)

Find the GCF of 63, 147, and 273
You can use prime factorisation to find the GCF of these integers.
63 = 3*3*7
147 = 3*7*7
273 = 3*7*13
Now select the factors that are common to the three lists of prime factors and multiply them.
These are 3 and 7
So the GCF = 3*7 = 21
Check:
63 = 3*21
147 = 7*21
273 = 13*21